# Symmetry-informed transferability of optimal parameters in the Quantum Approximate Optimization Algorithm

### v1

@@ Abstract
One of the main limitations of variational quantum algorithms is the classical optimization of the highly dimensional non-convex variational parameter landscape. To simplify this optimization, we can reduce the search space using problem symmetries and typical optimal parameters as initial points if they concentrate. In this article, we consider typical values of optimal parameters of the Quantum Approximate Optimization Algorithm for the MaxCut problem with $d$-regular tree subgraphs and reuse them in different graph instances. We prove symmetries in the optimization landscape of several kinds of weighted and unweighted graphs, which explains the existence of multiple sets of optimal parameters. However, we observe that not all optimal sets can be successfully transferred between problem instances. We find specific transferable domains in the search space and show how to translate an arbitrary set of optimal parameters into the adequate domain using the studied symmetries.
@@ I Introduction
Variational quantum algorithms (VQAs) are among the most popular and accessible quantum algorithms for solving classically hard problems, with a notable example being the Quantum Approximate Optimization Algorithm (QAOA) [1], used to tackle combinatorial optimization problems (COPs). The QAOA is a hybrid algorithm consisting of a quantum circuit with several layers of parameter-dependent operations applied iteratively, followed by a classical optimization of the parameters. The quality of the solutions output by this algorithm depends on these classically optimized variational parameters and the depth of the quantum circuit. In theory, the solution improves as the circuit depth increases but, in practice, deep circuits introduce noise that affects the quantum state. However, shallow depth QAOA is often inferior to state-of-the art classical algorithms [2, 3, 4, 5, 6]. Therefore, we have to choose an algorithm depth that achieves an acceptable solution while maintaining quantum coherence. Alternatively, one can modify the standard QAOA to mitigate the effects of decoherence. Multiple alternative QAOA inspired algorithms adapted to specific COPs show better theoretical performance at lower depths. Yet their success also depends on finding good variational parameters.
@@
Optimizing the variational parameters is an active field of research, and a multitude of techniques have been studied, including standard classical optimizers [14], gradient-free optimizers [14, 15], classical genetic algorithms [16], machine-learning assisted optimization [17, 18, 19], and extrapolation techniques [7, 20]. Furthermore, we can benefit from prior knowledge of the problem, either by initializing a quantum state close to the solution [21, 22, 23] or bypassing the optimization by transferring pre-trained optimal parameters [24, 25, 26, 27, 28, 29, 30, 31, 32]. The optimization can also be simplified by employing fixed schemes such as Fourier-series extrapolation [7], discrete adiabatic paths [7, 20], or linear ramp schedules [33]. While the main purpose of QAOA is finding approximate solutions to COPs, the sampling from the output quantum state is related to pseudo-Boltzmann distributions that are classically intractable [34, 35]. In these cases, the quality of the algorithm parameters affects the effective temperature of the probability distribution, with the optima leading to the lowest temperature.
In some cases, the optimal parameters concentrate on typical values in parameter space. This behavior was initially observed for MaxCut on $3$-regular graphs [36], and later for MaxCut on other kinds of graphs [7, 34, 37, 28]. This phenomenon has been linked with the locality of QAOA at shallow depths, where the cost Hamiltonian of a graph can be reformulated as a sum of local Hamiltonians of subgraphs [1, 7]. The typical optimal parameters are those optimal to the local Hamiltonian that dominate the expectation value of the total cost Hamiltonian [24, 26, 38, 25]. At algorithm depth $p=11$, numerical studies show that MaxCut on $3$-regular graphs with typical parameters has performance guarantees superior to the best classical algorithms [39]. Furthermore, it is possible to achieve near-optimal QAOA performances by transferring the typical values from a donor problem instance to a receiver problem instance drawn from a different ensemble. These successful transfers have been numerically confirmed between regular graphs with either even or odd degrees [25, 40], between the Sherrington-Kirkpatrick model and MaxCut on high-girth graphs [37, 36], between unweighted and weighted graphs [27, 28], and between the MaxCut problem and other COPs [29].
@@
In this article, we identify symmetries in the optimization landscape of QAOA for MaxCut that include certain kinds of weighted graphs, previously not addressed in the literature. In addition, we study the performance of QAOA using transferred parameters between problem instances for circuit depths $p=1$ and $p=2$. We define a function characterizing the transferability of parameters and provide an analytical closed-form expression for triangle-free regular graphs. We observe that, in general, these transferred parameters reach a high performance if they belong to certain domains in the search space. Our analysis extends to $d$-regular graphs with integer weights and unweighted random graphs from different models.
The article is structured as follows. In Sec. II, we review the formulation of QAOA and the COP MaxCut, as well as the known symmetries of the variational parameters. We prove a generalization of these symmetries to include a wider family of graphs. Then, in Sec. III, we analyze the transferability of optimal variational parameters and identify which optimal sets are suitable for reuse. Lastly, we summarize our results in Sec. IV.
@@ II Symmetries of QAOA for MaxCut
Table 1: Transformations of $(\bm{\gamma},\bm{\beta})\to(\bm{\gamma}^{\prime},\bm{\beta}^{\prime})$ such that $C(\bm{\gamma},\bm{\beta})=C(\bm{\gamma}^{\prime},\bm{\beta}^{\prime})$. The EWS symmetry refers to integer-weighted graphs where the sum of weights connected to each vertex is even, i.e., $\Sigma_{j}w_{jk}=2\mathbb{Z},\,\forall k\in V$. Analogously, OWS symmetry refers to integer-weighted graphs where the sum of the weights connected to each vertex is odd, i.e., $\Sigma_{j}w_{jk}=\mathbb{Z}\setminus 2\mathbb{Z},\,\forall k\in V$.
@@
The cost in Eq. (4) is invariant under parameter changes $\gamma_{i}\to\gamma_{i}\pm\pi$ for any layer $i$ in QAOA, for graphs where the sum of the weights connected to each vertex is even, i.e., $\Sigma_{j}w_{jk}=2\mathbb{Z},\forall k\in V$ (see proof in Appendix A). This even weight sum (EWS) symmetry captures the previously known result for unweighted graphs where all vertices have even degree [32]. More examples that exhibit this parameter pattern include graphs with only even weights and graphs where all vertices have even degrees with only odd weights, for instance, $w_{ij}\in\{-1,1\}$.
@@
The cost in QAOA is invariant under the simultaneous parameter change $\gamma_{i}\to\gamma_{i}\pm\pi$ and $\beta_{j}\to-\beta_{j}$ in all layers $j\geq i$, for graphs where the sum of the weights connected to each vertex is odd, i.e., $\Sigma_{j}w_{jk}=\mathbb{Z}\setminus 2\mathbb{Z},\forall k\in V$ (see proof in Appendix A). This odd weight sum (OWS) symmetry contains the case of unweighted graphs when all vertices have odd degree [32]. Moreover, we also include graphs where each vertex has an odd degree and all weights are odd, for example, $w_{ij}\in\{-1,1\}$.
@@ III Transferability of optimal parameters
The number of possible subgraphs for a $d$-regular graph grows exponentially with the algorithm depth $p$, if QAOA does not see the whole graph, i.e., if $p\sim\mathcal{O}(\log n)$ with $n$ the number of vertices. However, despite this exponential growth, in a sufficiently large random $d$-regular graph, almost all subgraphs will be the cycle-free tree subgraph [38, 2, 22]. Consequently, the optimal QAOA parameters for a $d$-regular graph will be close to the optimal parameters of the $d$-regular tree subgraph, which explains the successful parameter transferability between them at shallow depths. However, this argument fails to explain whether we can reuse optimal parameters between graphs with different degrees (or average degrees). Particularly, prior research suggests that, on average, we can successfully transfer any optimal parameters between regular graphs with the same parity (both have either odd or even degrees). However, for graphs with opposite parity, we can only reuse a subset of the optimal parameters [25, 40].
@@
with $k=C_{\text{max}}/|E|$, the fraction of edges cut in the maximum cut. In general, $k\in(1/2,1]$ and $k=1$ for bipartite graphs, and Eq. (7) is valid for sufficiently large graphs [1], that is, graphs of order larger than $n\sim(d-1)^{2p}$ (see Appendix B for a detailed derivation).
The function in Eq. (7) allows us to calculate the transferability error for regular graphs up to an arbitrary degree. We visualize this expression in Fig. 2 for bipartite receiver graphs with degrees up to $d=70$ and tree donor subgraphs of degree $d^{\prime}\in\{3,4,5\}$. The transferability error $\eta_{{d},{d^{\prime}}}^{g>3}$ is minimized for donor and receiver graphs with the same degree $d=d^{\prime}$, in agreement with previous results [24, 25]. We observe that as the difference between the receiver and donor degrees $d$ and $d^{\prime}$ grows, QAOA with the parameters transferred from the tree subgraphs performs as random guessing with an approximation ratio of $r=1/2$ [45]. Therefore, while the error tends to 00 as $d\to\infty$, transferring parameters becomes pointless in the large $d$ limit. Moreover, when the error eventually reaches 00 for large $d$, it also indicates that the approximation ratio of QAOA with depth $p=1$ for any optimal parameters is $r\approx 1/2$ [2].
@@ III.1 Numerical analysis
In this case, we use again the unweighted $d^{\prime}$-regular tree subgraph as a donor graph, for general random receiver graphs. Unlike regular graphs, random graphs do not have a fixed degree and, at low depths, QAOA sees a local subgraph with a varying degree around each vertex. The random graph models will have different distributions of degrees. Here, we study three models of random graphs: the Erdős-Rényi (ER) model [46], the Barabási-Albert (BA) model [47], and the Watts-Strogatz (WS) model [48].
@@
As we show in Sec. II, at depth $p=2$, QAOA for MaxCut on $d$-regular graphs has eight sets of optimal parameters in $\mathcal{A}^{2}$, instead of the four sets at $p=1$. We present these parameters in Appendix C (see Table 2) for the $3$-regular tree subgraph. Here, we study the performance of QAOA for unweighted regular graphs of order $n=120$ and degrees $d\in\{3,4,5\}$, considering ten graphs per degree, and using—without shifting to a common domain—the eight sets of optimal parameters of the $3$-regular tree donor subgraph.
@@
IV Conclusions and outlook
Optimizing the variational circuits remains a big obstacle for realizable and successful VQAs. One useful simplification in the optimization task is limiting the search space using symmetries in the energy landscape. Additionally, the optimal parameters concentrate on typical values for different instances of certain problems. Thus, we can use these typical values as pretrained parameters and transfer them between instances, further simplifying the optimization process.

### v2

@@ Abstract
One of the main limitations of variational quantum algorithms is the classical optimization of the highly dimensional non-convex variational parameter landscape. To simplify this optimization, we can reduce the search space using problem symmetries and typical optimal parameters as initial points if they concentrate. In this article, we consider typical values of optimal parameters of the quantum approximate optimization algorithm for the MaxCut problem with $d$-regular tree subgraphs and reuse them in different graph instances. We prove symmetries in the optimization landscape of several kinds of weighted and unweighted graphs, which explains the existence of multiple sets of optimal parameters. However, we observe that not all optimal sets can be successfully transferred between problem instances. We find specific transferable domains in the search space and show how to translate an arbitrary set of optimal parameters into the adequate domain using the studied symmetries. Finally, we extend these results to general classical optimization problems described by Ising Hamiltonians, the Hamiltonian variational ansatz for relevant physical models, and the recursive and multi-angle quantum approximate optimization algorithms.
@@ I Introduction
Variational quantum algorithms (VQAs) are among the most popular and accessible quantum algorithms for solving classically hard problems, with a notable example being the quantum approximate optimization algorithm (QAOA) [1], used to tackle combinatorial optimization problems (COPs). The QAOA is a hybrid algorithm consisting of a quantum circuit with several layers of parameter-dependent operations applied iteratively, followed by a classical optimization of the parameters. The quality of the solutions output by this algorithm depends on these classically optimized variational parameters and the depth of the quantum circuit. In theory, the solution improves as the circuit depth increases but, in practice, deep circuits introduce noise that affects the quantum state. However, shallow depth QAOA is often inferior to state-of-the art classical algorithms [2, 3, 4, 5, 6]. Therefore, we have to choose an algorithm depth that achieves an acceptable solution while maintaining quantum coherence. Alternatively, one can modify the standard QAOA to mitigate the effects of decoherence. Multiple alternative QAOA inspired algorithms adapted to specific COPs show better theoretical performance at lower depths, like the recursive QAOA (RQAOA) [7] or the multi-angle QAOA (ma-QAOA) [8]. Yet their success also depends on finding good variational parameters.
@@
Optimizing the variational parameters is an active field of research, and a multitude of techniques have been studied, including standard classical optimizers [16], gradient-free optimizers [16, 17], classical genetic algorithms [18], machine-learning assisted optimization [19, 20, 21, 22], and extrapolation techniques [9, 23]. Furthermore, we can benefit from prior knowledge of the problem, either by initializing a quantum state close to the solution [24, 25, 26], finding a better initial ansatz by an efficient classical optimization over Clifford gates [27, 28] or bypassing the optimization by transferring pre-trained optimal parameters [29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. The optimization can also be simplified by employing fixed schemes such as Fourier-series extrapolation [9], discrete adiabatic paths [9, 23], or linear ramp schedules [39]. While the main purpose of QAOA is finding approximate solutions to COPs, the sampling from the output quantum state is related to pseudo-Boltzmann distributions that are classically intractable [40, 41]. In these cases, the quality of the algorithm parameters affects the effective temperature of the probability distribution, with the optima leading to the lowest temperature.
In some cases, the optimal parameters concentrate on typical values in parameter space. This behavior was initially observed for MaxCut on $3$-regular graphs [42], i.e., graphs where every vertex is connected to 3 vertices, and later for MaxCut on other kinds of graphs [9, 40, 43, 33]. This phenomenon has been linked with the locality of QAOA at shallow depths, where the cost Hamiltonian of a graph can be reformulated as a sum of local Hamiltonians of subgraphs [1, 9]. The typical optimal parameters are those optimal to the local Hamiltonian that dominate the expectation value of the total cost Hamiltonian [29, 31, 44, 30]. At algorithm depth $p=11$, numerical studies show that MaxCut on $3$-regular graphs with typical parameters has performance guarantees superior to the best classical algorithms [45], and a recent work by Ref. [46] suggests that the QAOA with typical parameters already has an advantage at $p=8$. Furthermore, it is possible to achieve near-optimal QAOA performances by transferring the typical values from a donor problem instance to a receiver problem instance drawn from a different ensemble. These successful transfers have been numerically confirmed between regular graphs with either even or odd degrees [30, 47], between the Sherrington-Kirkpatrick model and MaxCut on high-girth graphs [43, 42], between unweighted and weighted graphs [32, 33], and between the MaxCut problem and other COPs [34]. Additionally, the transferred parameters can be used as initial guesses for a warm-start local optimization of QAOA parameters, significantly accelerating the classical loop [19, 9, 48, 35]. Warm-start optimization is a promising heuristic method to mitigate the challenge of barren plateaus. Unlike other methods to avoid barren plateaus which often inadvertently imply that the problem at hand is efficiently solvable by classical computers, warm-start techniques are not believed to have this limitation [14].
@@
In this article, we identify symmetries in the optimization landscape of QAOA for Ising Hamiltonian with certain integer coefficients previously not addressed in the literature. In addition, we study the performance of QAOA for MaxCut using transferred parameters between problem instances for circuit depths $p=1$ and $p=2$. We define a function characterizing the transferability of parameters and provide an analytical closed-form expression for triangle-free regular graphs. Such graphs are said to have a girth higher than 3, i.e., the shortest cycle length is longer than 3. We observe that, in general, these transferred parameters reach a high performance if they belong to certain domains in the search space. Our analysis extends to $d$-regular graphs with integer weights and unweighted random graphs from different models—graphs with a fixed number of vertices, and edges generated by different probability distributions. We show that when specific parameters are transferable, all parameters can be effectively transferred by applying translations based on the symmetries of the variational state’s expectation value. This result appears to extend to higher algorithm depths, suggesting a scalable approach to parameter optimization. Furthermore, while our numerical study focuses on particular problem instances, our work provides tools to analyze the symmetries and transferability of parameters, which can help to identify successful initialization or optimization schemes in general VQAs. Building on this idea, we extend our symmetry results to the Hamiltonian variational ansatz (HVA), a relevant quantum heuristic for physical models in the context of variational quantum eigensolvers (VQEs), as well as the RQAOA and the ma-QAOA. For the latter two, we also discuss the potential of parameter transferability.
Different initialization or optimization schemes have unique strengths and weaknesses that make them more suitable for specific VQAs. For hardware-efficient ansätze, a classically efficient search over Clifford-gate constructed ansätze has proven effective for finding good initializations [28]. However, this approach is less suitable for structured ansätze that utilize problem-specific features, such as QAOA or HVA circuits. Indeed, in QAOA circuits, Clifford gates constrain the parameters to integer multiples of $\pi/4$, which are typically far from the optimal values [27], unless one considers hardware-efficient QAOA variants, like ADAPT-QAOA [50]. Conversely, our results are not directly applicable to hardware-efficient ansatz circuits, as we rely on integer-valued coupling constants in the spin model to derive parameter space symmetries. These assumptions suggest that our findings will improve parameter training in structured quantum circuits for specific problems with integer-valued coupling constants in the Hamiltonians.
The article is structured as follows. In Sec. II, we review the formulation of QAOA and the COP MaxCut, as well as the known symmetries of the variational parameters. We prove a generalization of these symmetries to include a wider family of graphs. Then, in Sec. III, we analyze the transferability of optimal variational parameters and identify which optimal sets are suitable for reuse. In Sec. IV, we generalize our symmetries to encompass more VQAs and discuss the parameters’ transferability in those cases. Lastly, we summarize our results in Sec. V.
@@ II Symmetries of QAOA for MaxCut
Table 1: Transformations of $(\bm{\gamma},\bm{\beta})\to(\bm{\gamma}^{\prime},\bm{\beta}^{\prime})$ such that $C(\bm{\gamma},\bm{\beta})=C(\bm{\gamma}^{\prime},\bm{\beta}^{\prime})$. The EWS symmetry refers to integer-weighted graphs where the sum of weights connected to each vertex is even, i.e., $\Sigma_{j}w_{jk}\in 2\mathbb{Z},\,\forall k\in V$. Analogously, OWS symmetry refers to integer-weighted graphs where the sum of the weights connected to each vertex is odd, i.e., $\Sigma_{j}w_{jk}\in\mathbb{Z}\setminus 2\mathbb{Z},\,\forall k\in V$.
@@
The cost in Eq. (4) is invariant under parameter changes $\gamma_{i}\to\gamma_{i}\pm\pi$ for any layer $i$ in QAOA, for graphs where the sum of the weights connected to each vertex is even, i.e., $\Sigma_{j}w_{jk}\in 2\mathbb{Z},\forall k\in V$ (see proof in Appendix B). This even weight sum (EWS) symmetry captures the previously known result for unweighted graphs where all vertices have even degree [37]. More examples that exhibit this parameter pattern include graphs with only even weights and graphs where all vertices have even degrees with only odd weights, for instance, $w_{ij}\in\{-1,1\}$.
@@
The cost in QAOA is invariant under the simultaneous parameter change $\gamma_{i}\to\gamma_{i}\pm\pi$ and $\beta_{j}\to-\beta_{j}$ in all layers $j\geq i$, for graphs where the sum of the weights connected to each vertex is odd, i.e., $\Sigma_{j}w_{jk}\in\mathbb{Z}\setminus 2\mathbb{Z},\forall k\in V$ (see proof in Appendix B). This odd weight sum (OWS) symmetry contains the case of unweighted graphs when all vertices have odd degree [37]. Moreover, we also include graphs where each vertex has an odd degree and all weights are odd, for example, $w_{ij}\in\{-1,1\}$.
@@ III Transferability of optimal parameters
The number of possible subgraphs for a $d$-regular graph grows exponentially with the algorithm depth $p$, if QAOA does not see the whole graph, i.e., if $p\sim\mathcal{O}(\log n)$ with $n$ the number of vertices. However, despite this exponential growth, in a sufficiently large random $d$-regular graph, almost all subgraphs will be tree subgraphs [44, 2, 25], i.e., connected subgraphs graphs with no cycles. Consequently, the optimal QAOA parameters for a $d$-regular graph will be close to the optimal parameters of the $d$-regular tree subgraph, which explains the successful parameter transferability between them at shallow depths. However, this argument fails to explain whether we can reuse optimal parameters between graphs with different degrees (or average degrees). Particularly, prior research suggests that, on average, we can successfully transfer any optimal parameters between regular graphs with the same parity (both have either odd or even degrees). However, for graphs with opposite parity, we can only reuse a subset of the optimal parameters [30, 47].
@@
with $k=C_{\text{max}}/|E|$, the fraction of edges cut in the maximum cut. In general, $k\in(1/2,1]$ and $k=1$ for bipartite graphs, and Eq. (7) is valid for sufficiently large graphs [1], that is, graphs of order larger than $n\sim(d-1)^{2p}$ (see Appendix E for a detailed derivation).
The function in Eq. (7) allows us to calculate the transferability error for regular graphs up to an arbitrary degree. We visualize this expression in Fig. 2 for bipartite receiver graphs with degrees up to $d=70$ and tree donor subgraphs of degree $d^{\prime}\in\{3,4,5\}$. We also include a reference line where QAOA with donated parameters is equivalent to random guessing, that is, when QAOA gives an approximation ratio of $\frac{1}{2}$, or equivalently $\eta_{{R},{D}}=r_{R}(\bm{\gamma}_{R}^{*},\bm{\beta}_{R}^{*})-\frac{1}{2}$. Note that if we shift this line by $\frac{1}{2}$, it corresponds to the approximation ratio of QAOA with optimal parameters, $r_{R}(\bm{\gamma}_{R}^{*},\bm{\beta}_{R}^{*})$, for increasing graph degrees. The transferability error $\eta_{{d},{d^{\prime}}}^{g>3}$ is minimized for donor and receiver graphs with the same degree $d=d^{\prime}$, in agreement with previous results [29, 30]. We observe that as the difference between the receiver and donor degrees $d$ and $d^{\prime}$ grows, QAOA with the parameters transferred from the tree subgraphs performs as random guessing with an approximation ratio of $r=1/2$ [53]. Therefore, while the error tends to 00 as $d\to\infty$, transferring parameters becomes pointless in the large $d$ limit. Moreover, when the error eventually reaches 00 for large $d$, it also indicates that the approximation ratio of QAOA with depth $p=1$ for any optimal parameters is $r\approx 1/2$ [2].
@@ III.1 Numerical analysis
In this case, we use again the unweighted $d^{\prime}$-regular tree subgraph as a donor graph, for general random receiver graphs. Unlike regular graphs, random graphs do not have a fixed degree and, at low depths, QAOA sees a local subgraph with a varying degree around each vertex. The random graph models will have different distributions of degrees. Here, we study three models of random graphs: the Erdős-Rényi (ER) model [54], the Barabási-Albert (BA) model [55], and the Watts-Strogatz (WS) model [56]. While these models can be tuned to have the same average degree, they exhibit distinct properties that make them suitable for describing different systems. The ER model generates random graphs where edges appear independently with equal probability, resulting in a binomial degree distribution. In contrast, the BA model generates scale-free networks characterized by large hubs, where a few vertices have many connections and most have few, following a power-law degree distribution. The WS model produces graphs with local clusters of connected vertices by rewiring edges in a ring lattice with fixed probability. Both BA and WS graphs are representative of real-world network structures.
@@
As we show in Sec. II, at depth $p=2$, QAOA for MaxCut on $d$-regular graphs has eight sets of optimal parameters in $\mathcal{A}^{2}$, instead of the four sets at $p=1$. We present these parameters in Appendix F (see Table 2) for the $3$-regular tree subgraph. Here, we study the performance of QAOA for unweighted regular graphs of order $n=120$ and degrees $d\in\{3,4,5\}$, considering ten graphs per degree, and using—without shifting to a common domain—the eight sets of optimal parameters of the $3$-regular tree donor subgraph.
@@
IV Parameter symmetries in other variational quantum algorithms
The quantum circuit of many hybrid classical-quantum algorithms can be built through several iterations of similar quantum operations, e.g., any layer of QAOA contains a parametrized mixing unitary operation and a parametrized cost unitary operation. Such structured quantum circuit ansätze also appear in realizations of VQEs for finding ground states of quantum many-body problems, as the HVA. Compared to the hardware-efficient ansatz, the problem-specific HVA circuit provides an easier optimization, reducing the challenges of barren plateaus while maintaining a high expressivity [57]. While in this paper our focus mainly lies on the QAOA for different MaxCut instances and the transferability of its optimal parameters between various ensembles of graphs, we discuss here how these results can be extended to other COPs and VQAs. The EWS and OWS symmetries shown in Sec. II for the MaxCut problem also apply to all Ising Hamiltonians of the form $\hat{H}_{C}=\sum_{i<j}w_{ij}\hat{\sigma}_{i}^{z}\hat{\sigma}_{j}^{z}/2+\sum_{i}h_{i}\hat{\sigma}_{z}^{i}/2$ (see detailed derivation in Appendix C). The condition on the coefficients $w_{ij}$ and $h_{i}$ change to $\sum_{j}w_{jk}+h_{k}\in 2\mathbb{Z},\forall k\in V$ and $\sum_{j}w_{jk}+h_{k}\in\mathbb{Z}\setminus 2\mathbb{Z},\forall k\in V$ for EWS and OWS, respectively. This extension covers a wide range of problems, including Karp’s 21 NP-complete problems [58] and finding ground states of spin glass models.
The quantum unitary operations of the HVA use terms of many-body Hamiltonians, analogously to the cost function in QAOA and adiabatic quantum computation. Therefore, we can identify similar symmetries in these VQAs. A notable example is the HVA for the transverse-field Ising model (TFIM), which coincides with the QAOA for solving MaxCut in the “ring of disagrees”, a $2$-regular unweighted ring graph [1]. Therefore, the HVA for the TFIM exhibits EWS, and the parameters can be optimized over a smaller domain. In Appendix D, we show that EWS is valid for ansätze constructed with operators of the form
$\hat{U}_{\alpha\alpha}(\gamma)=\prod_{(i,j)\in E}\exp(-i\frac{w_{ij}}{2}\gamma\hat{\sigma}_{\alpha}^{i}\hat{\sigma}_{\alpha}^{j}),$ (8)
where $\alpha\in\{x,y,z\}$, given that the weights $w_{ij}$ fulfil the criteria of EWS described in Sec. II. In other words, any VQE ansätz with these terms and conditions follows the EWS.
We extend our results to other HVA based on different Hamiltonian terms. In particular, we consider the general case of a Heisenberg model on high-dimensional spaces, where the connectivity between the spins is defined by the set of edges $E$ of a given graph. Moreover, we allow the coupling constants between the spins to vary from site to site, thus introducing disorder. The Hamiltonian reads
$\hat{H}=\frac{1}{2}\sum_{(i,j)\in E}\left(w_{ij}^{x}\hat{\sigma}_{x}^{i}\hat{\sigma}_{x}^{j}+w_{ij}^{y}\hat{\sigma}_{y}^{i}\hat{\sigma}_{y}^{j}+w_{ij}^{z}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{j}\right),$ (9)
with $w_{ij}^{x}$, $w_{ij}^{y}$, and $w_{ij}^{z}$ the coupling strengths associated with the $x$-, $y$-, and $z$-components of the spin interaction between the sites $i$ and $j$, given by the edge $(i,j)$. This general case reduces to well-known models for studying magnetic systems. For instance, we can consider a quantum spin chain with real-valued coupling constants, such that the Hamiltonian becomes
$\hat{H}_{\rm{XYZ}}=\sum_{i=1}^{n}\left(J_{x}\hat{\sigma}_{x}^{i}\hat{\sigma}_{x}^{i+1}+J_{y}\hat{\sigma}_{y}^{i}\hat{\sigma}_{y}^{i+1}+J_{z}\hat{\sigma}_{z}^{i}\hat{\sigma}_{z}^{i+1}\right),$ (10)
in which we assume periodic boundary conditions ($11\equiv n+$ 1). This one-dimensional Heisenberg XYZ model corresponds to a particular case of Eq. (9) for a $2$-regular ring graph with $n$ vertices defining the connectivity of the interactions, and uniform coupling constants $w_{ij}^{\alpha}/2=J_{\alpha}$, with $\alpha\in\{x,y,z\}$. When $J_{x}=J_{y}$, the model is called the one-dimensional Heisenberg XXZ model, studied in the context of HVA [57]. Without loss of generality, we can take $J_{x}=J_{y}=1$ and $J_{z}=\Delta$, with $\Delta$ controlling the spin anisotropy. The critical point $\Delta=1$ becomes harder to approximate classically [57]. Interestingly, optimal HVA parameters have been successfully transferred from small to large instances of the Heisenberg XXZ model for parameters in the domain $\mathcal{U}_{1}^{\,p}$ [38]. We derive symmetry results for all these models in the case of integer-valued coupling constants $w_{ij}^{\alpha}$. In Appendix D, we base our derivations on the fact that the HVA for the general Hamiltonian of Eq. (9) comprises products of operators given by Eq. (8).
Finally, we also consider different versions of the QAOA, namely the RQAOA [7] and the ma-QAOA [8]. The RQAOA reduces the original optimization problem size by imposing constraints obtained from recursive applications of QAOA. With each algorithm iteration, the weights and regularity of the problem graph change, altering the cost Hamiltonian for the QAOA subroutine. As we show in Appendix D, RQAOA preserves the EWS of the original cost Hamiltonian throughout every iteration—each reduced cost Hamiltonian adheres to the EWS—but breaks the OWS after the first iteration. This behavior implies that all optimal parameters are directly transferable if the original graph obeys EWS. Otherwise, for initial problem graphs with OWS, we can see from our discussion on transferability to random model graphs in Sec. III, that only optimal parameters from the domains $\mathcal{U}_{1}$ or $\mathcal{U}_{2}$ can be reused between iterations. Another version of QAOA is the ma-QAOA, which uses a higher number of parameters at shallow depths to create a more expressive ansatz. Unlike standard QAOA, which requires EWS or OWS for the entire graph, with ma-QAOA we can focus on the weight sums of individual vertices to analyze the transferability of the gate parameters related to Hamiltonian terms of specific vertices and edges. Both RQAOA and ma-QAOA improve the performance of the standard QAOA at shallow depths, albeit with increased demands on parameter optimization [8, 59, 7, 60]. Therefore, insights from transferability analyses to improve these optimization loops are especially valuable.
V Conclusions and outlook